A different constant velocity cart travelled 0.980.01 m in 4.20.1s.

Calculate the speed that the car traveled, including uncertainty. Show and explain ALL your working.
1
Expert's answer
2020-05-08T16:06:35-0400

The speed:


"v=\frac{s}{t}=\frac{0.98}{4.2}=0.233\frac{m}{s}"

Uncertainty:


"\frac{\Delta v}{v}=\frac{\Delta s}{s}+\frac{\Delta t}{t}=\frac{0.01}{0.98}+\frac{0.1}{4.2}=0.034"

"\Delta v=(0.034)(0.23)=0.008\frac{m}{s}"